OpenStudy (anonymous):

this is a question in topology, prove that if A is a countable dense set of IR then there exists a subset B of A dense and countable as well.

OpenStudy (ikram002p):

hey :D

OpenStudy (ikram002p):

so would you mind to write countable dense def ? tbh this proof is rigid xD

OpenStudy (anonymous):

Hi, I'll start first with def of a dense set we say that A is dense if $\bar{A}=\mathbb{R}$

OpenStudy (anonymous):

yes sorry

OpenStudy (ikram002p):

u ment this , right ? \( \bar{A}=\mathbb{R} \)

OpenStudy (anonymous):

\[\bar{A}=\mathbb{R}\]

OpenStudy (ikram002p):

ohk :D continue

OpenStudy (anonymous):

and A is countable if there is an injective function from A to \[\mathbb{N}\]

OpenStudy (ikram002p):

well do you know that every infinite set has a countable dense subset ?

OpenStudy (anonymous):

no actually that is a part of what we're trying to prove , no?

OpenStudy (ikram002p):

well i was trying to reach the proof by showing that there exist A infinite group ... mmm but since you dnt know it , lets try something else

OpenStudy (anonymous):

okay

OpenStudy (ikram002p):

so since A is a countable dense set of IR then 1-\(\bar{A}=\mathbb{R}\) 2- there is F(x)=y , s.t f(x) is 1_1 (injective ) i still cant catch a line -.- i mean what i know is each subset of sense set is also dense also , if a set have 1_1 injective function then there exist a subset that also have 1_1 injective .. im lost with this , sry xD

OpenStudy (anonymous):

it's okay that made me confused too :p

OpenStudy (ikram002p):

:D